OpenStudy (anonymous):

a magnetic flux through a stationary loop with a resistance R varies v as l=at(v-t). find the amount of heat generated in the loop during that time . The inductance of the loop is to be neglected .

OpenStudy (anonymous):

H = I^2 x R x t I =V/R V=d(flux)/dt